### v1

# Distribution Shifts at Scale: Out-of-distribution Detection in Earth Observation

Abstract
Training robust deep learning models is critical in Earth Observation, where globally deployed models often face distribution shifts that degrade performance, especially in low-data regions. Out-of-distribution (OOD) detection addresses this challenge by identifying inputs that differ from in-distribution (ID) data. However, existing methods either assume access to OOD data or compromise primary task performance, making them unsuitable for real-world deployment. We propose TARDIS, a post-hoc OOD detection method for scalable geospatial deployments. The core novelty lies in generating surrogate labels by integrating information from ID data and unknown distributions, enabling OOD detection at scale. Our method takes a pre-trained model, ID data, and WILD samples, disentangling the latter into surrogate ID and surrogate OOD labels based on internal activations, and fits a binary classifier as an OOD detector. We validate TARDIS on EuroSAT and xBD datasets, across 17 experimental setups covering covariate and semantic shifts, showing that it performs close to the theoretical upper bound in assigning surrogate ID and OOD samples in 13 cases. To demonstrate scalability, we deploy TARDIS on the Fields of the World dataset, offering actionable insights into pre-trained model behavior for large-scale deployments. The code is publicly available at https://github.com/microsoft/geospatial-ood-detection
Figure 1: Illustration of the proposed method for OOD detection. Given a pre-trained model, a dataset of known ID samples, and WILD samples (containing both ID and OOD data), the method assigns surrogate ID or OOD labels to each WILD sample and fits a binary classifier $g$ (top row). During deployment, the classifier $g$ uses internal activations to predict whether an unseen sample is ID or OOD (bottom row).
1 Introduction
Deep learning (DL) models have demonstrated remarkable capabilities across various domains but often exhibit overconfidence in their predictions, even when confronted with data that diverges from their training distribution [25, 10, 19]. This overconfidence arises from the assumption that inference data will follow the same independent and identically distributed (i.i.d.) properties as the training data. However, in real-world applications, this closed-world assumption [8, 14] is frequently violated by test-time distribution shifts (e.g., lighting or angle variations, different devices or sensors) that can significantly degrade model performance and harm generalization. To address this, it is critical for predictive models to detect when new observations fall outside the training distribution, a task known as out-of-distribution (OOD) detection. Differentiating OOD from in-distribution (ID) samples is particularly challenging, partly due to the poor calibration of neural networks [6, 18]. This degradation poses a critical challenge, especially in applications where incorrect predictions can have severe consequences. In Earth Observation (EO), specifically with satellite imagery, distribution shifts are even more common due to variability in sensor types, acquisition angles, temporal changes, spatial variations, and landscape alterations caused by events such as natural disasters. This variability necessitates unique, tailored approaches for satellite imagery [21]. The nature of distribution shifts between training and inference data can be classified as near-distribution or far-distribution shifts, each posing unique challenges. For example, a model trained on data from a certain season and region may perform poorly when applied to other seasons in the same region or the same season in different regions [13], representing a far-distribution shift. Even models trained on data covering multiple seasons and regions may perform suboptimally under specific acquisition or atmospheric conditions, such as cloud cover, sensor faults, or artifacts, indicative of a near-distribution shift.
Figure 2: The proposed framework consists of four key steps: (1) Sampling in-distribution (ID) and WILD samples; (2) Extracting internal activations from a pre-trained model $f$ for both ID and WILD samples; (3) Clustering the combined feature space and labeling WILD samples as surrogate-ID or surrogate-OOD; (4) Fitting a binary classifier $g$ on the labeled feature representations to distinguish between ID and OOD samples. The classifier $g$, during deployment, flags out-of-distribution inputs.
Despite the critical importance of robust models in Earth Observation (EO), only a few studies address OOD detection in this domain [5, 3, 1]. These studies typically assume access to test-time distributions, rely on closed-set assumptions, and often sacrifice model performance on in-distribution tasks. This highlights the need for tailored OOD detection methods suited for large-scale EO deployment.
A further challenge arises when applying geospatial models globally. Many geospatial products, while measuring the same ground realities with high coverage, exhibit low levels of agreement [2]. Additionally, global geospatial models often suffer significant performance drops in low-data regions, such as sub-Saharan Africa [12]. These issues emphasize the need for global models equipped to detect OOD samples during inference.
OOD detection capabilities in real-world applications enhance model reliability by identifying when input data deviates from the training distribution. They also support resource allocation and targeted data collection by pinpointing areas where the model may underperform, effectively guiding both computational and human resources. Additionally, OOD detection enables real-time adjustments, allowing models to remain reliable in dynamic conditions. This approach supports a human-in-the-loop strategy for critical decisions, fostering trust, transparency, and accountability. Overall, OOD detection is essential for managing large-scale EO deployments effectively.
In this paper, we begin with an essential observation: OOD samples can trigger internal activation patterns that diverge from those of ID samples. Building on this insight and recognizing the challenges of deploying geospatial models at scale, we introduce the following key contributions:
  • We propose TARDIS (Test-time Addressing of Distribution Shifts at Scale), a post-hoc OOD detection method that preserves the model’s performance on the main task and operates without requiring access to unknown test-time near-distribution shifts.
  • We evaluate TARDIS on EuroSAT, a patch-classification dataset, and xBD, a semantic segmentation dataset, examining both covariate and semantic shifts across 16 experimental setups. For covariate shifts, we introduce geographical variance (spatial changes), temporal variance (seasonal changes), and environmental variance (landcover or disaster type changes). Semantic shifts are introduced by withholding a class during training and exclusively testing on that class, thus presenting new semantic categories.
  • We demonstrate that TARDIS can be deployed at scale, providing actionable insights into the trustworthiness and robustness of geospatial models in real-world scenarios.
2 Related Works
Our method lies at the intersection of two notable categories of OOD detection methods: scoring functions and activation manipulation. Scoring functions assign a numerical score to each input, reflecting its alignment with ID samples based on the model’s output. Maximum Softmax Probability (MSP) [10] detects OOD samples by assessing the softmax confidence score, assuming that low confidence indicates OOD samples. The Out-of-DIstribution detector for Neural networks (ODIN) [16] enhances MSP by applying input perturbations and temperature scaling to improve ID-OOD separation. The Mahalanobis score [15] calculates the distance between input features and class means in feature space, flagging inputs far from these means as OOD. The energy score [17] evaluates the model’s energy function to assess the likelihood of an input belonging to the ID distribution. Another category of OOD detection methods manipulates the internal activations of a pre-trained model to improve detection performance. ReAct [24] identifies differences in activation patterns of the penultimate layer between ID and OOD samples, enhancing separation by clipping activations at an upper limit. DICE [23] applies weight sparsification on a specific layer to further distinguish ID from OOD data and, when combined with ReAct, can enhance detection performance. Activation Shaping [4] prunes a portion of an input sample’s activations and slightly adjusts the remaining activations; when paired with the energy score, this approach has been shown to outperform contemporary OOD methods. Neuron Activation Patterns (NAP) [20] extracts, downsamples, and binarizes activation patterns from convolutional layers, computing the smallest Hamming distance between binarized test and training patterns. This distance provides a measure of model uncertainty, aiding in OOD detection.
Both scoring functions and activation-based strategies face notable challenges in dynamic, real-world environments. Scoring functions, while maintaining ID accuracy, struggle with OOD detection due to their reliance on static data distributions, an assumption rarely valid in practice. Activation-based methods, though effective, often compromise ID task performance and rely on hyperparameters tuned with hypothetical or substitute OOD samples. This reliance limits their effectiveness under evolving test-time conditions and shifting distributions.
The task of OOD detection is inherently challenging, and the EO domain is no exception due to the diverse and heterogeneous nature of satellite imagery. Despite the critical importance of OOD detection in EO, this area remains relatively underexplored, with only a limited number of studies addressing the issue. One such study develops a Dirichlet Prior Network to quantify distributional uncertainty in DL models for satellite image analysis [5]. This method assumes Dirichlet distributions for ID samples and employs various setups where classes, color channels, and environmental features are alternately treated as ID or OOD. Another study frames anomaly detection as a cumulative open-set detection and location separation task [3]. This approach also uses a Dirichlet prior, similar to the first study, to expand the representation space between ID (normal) and OOD (anomalous) samples by predicting the anticipated categorical distribution. The method is suitable for scenarios where pre-event images are either unavailable, subject to radiation differences, or not sufficiently recent to aid in detection.
In this paper, we do not aim to optimize OOD detection performance but instead propose a method designed for real-world applications where the distribution is unknown, maintaining ID task performance is critical, and computationally expensive methods are impractical. Our focus is on handling near-distribution shifts, where data variations stem from the same satellite sources but exhibit changes over time, geography, or environmental factors. This contrasts with far-distribution shifts, where variations arise from entirely different data sources, such as different satellite models or natural images.
3 Problem Formulation
Let $\mathcal{X}$ represent the set of all possible data the model may encounter, and $\mathcal{Y}$ the set of class labels. The dataset on which a model is trained is defined as in-distribution (ID), denoted $D_{\text{ID}}\subset\mathcal{X}_{\text{ID}}\times\mathcal{Y}$. During inference, however, the model may encounter data from an unknown distribution, referred to as the WILD dataset $D_{\text{WILD}}\subset\mathcal{X}_{\text{WILD}}$, which may contain both ID and OOD samples.
We assume a pre-trained neural network $f:\mathcal{X}\rightarrow\mathbb{R}^{|\mathcal{Y}|}$, trained on $D_{\text{ID}}$, and our objective is to distinguish between ID and OOD samples within $D_{\text{WILD}}$. The network $f$ extracts features $\mathbf{z}\in\mathbb{R}^{F}$, where $F$ is the dimensionality of the feature space. Specifically, we fit a binary classifier $g:\mathbb{R}^{F}\rightarrow\{0,1\}$, parametrized by $\theta$, that operates on these features $\mathbf{z}$ for a given sample in $\mathcal{X}_{\text{WILD}}$. We define $g(\mathbf{z};\theta)=0$ if $\mathbf{x}\sim D_{\text{ID}}$, and $g(\mathbf{z};\theta)=1$ if $\mathbf{x}\sim D_{\text{OOD}}$.
4 Method
Fitting a classifier to distinguish between in-distribution (ID) and out-of-distribution (OOD) samples typically requires known ID and OOD labels, which is impractical in real-world deployments. This creates a need for surrogate ID and OOD labels for samples with unknown distributions, which we refer to as WILD samples. Our method, illustrated in Figure 1, operates with a given model, a set of known ID samples, and a set of WILD samples. It generates surrogate ID and OOD labels for the WILD samples, which are then used to train a binary classifier that can classify the distribution of incoming samples.
First, we sample $M$ in-distribution samples from the ID set and $N$ WILD samples from the WILD set. Next, we pass both the ID and WILD samples through the frozen pre-trained model and extract activations from a specified layer. A pooling operation is applied to perform spatial downsampling, transforming the activations into one-dimensional vectors and combining the downsampled features. We then use k-means clustering to partition the combined features into $K$ clusters. For each cluster, we calculate the proportion of ID samples and assign surrogate labels based on a threshold $T$: if the proportion of ID samples in the cluster is greater than or equal to $T$, the surrogate label is 0; otherwise, it is 1.
Here, $y_{\text{sur}}(\mathbf{z})$ denotes the surrogate label: 00 for surrogate ID and $1$ for surrogate OOD. With this labeled dataset, we fit a binary classifier $g(\mathbf{z};\theta)$, which maps a sample to a predicted label. During deployment, for a sample of unknown distribution, we perform activation extraction and classification as follows: $\hat{y}=g\left(\text{Downsample}\left(f(\mathbf{x})\right);\theta\right)$. The output can be interpreted as probability scores, allowing it to quantify the magnitude of domain shift or thresholded to produce binary values (00 for in-distribution and $1$ for out-of-distribution).
Surrogate Label Assignment. The core novelty of TARDIS lies in its surrogate labeling generation step, which enables us to assign labels to WILD samples based on their proximity to ID samples in feature space. The assumption is that ID and OOD features are sufficiently distinct to allow effective separation through clustering. Specifically, we use k-means clustering to label WILD samples as surrogate OOD if WILD samples outnumber ID samples in a cluster, and as surrogate ID otherwise. Without access to true OOD data, ID samples serve as an indirect benchmark for distinguishing between ID and WILD samples. To optimize this process, we tune the number of clusters $K$ and threshold $T$ by balancing three objectives: entropy (to encourage homogeneity within clusters and distinct groups of surrogate labels), correct ID proportion (to maximize accurate identification of ID samples), and incorrect ID proportion (to minimize the misclassification of ID samples as OOD). The combined objective function minimizes average entropy and the misclassification rate of ID samples while maximizing correct ID classification, resulting in well-separated and accurately labeled groups.
Figure 2 illustrates our method. First, ID samples are drawn from the model’s training set, while WILD samples, representing unknown distributions, are sourced from satellite images taken at any location and time (Step 1). Both ID and WILD samples are passed through a pre-trained model, and internal activations are extracted from a specified layer (Step 2). Using these activations, we assign surrogate labels to WILD samples, categorizing them as surrogate ID or surrogate OOD (Step 3). This surrogate labeling, central to our approach, decomposes WILD samples based on their distributional alignment with ID samples. Finally, these labeled feature representations are used to train a binary classifier that, during deployment, outputs probability scores or thresholded 0/1 predictions to classify samples as either ID or OOD (Step 4).
Table 1: Overview of experimental setups for evaluating TARDIS under controlled distribution shifts. The table details the covariate and semantic shift scenarios, including the dataset used, split method, and the composition of the ID and out-of-distribution OOD sets. All conditions are post-event except where specified.
Type
Dataset
Split Method ID (Training Set) OOD (Test Set)
Covariate Shift
EuroSAT
Spatial Split (Longitude) West Europe East Europe
xBD Similar Disasters - Distant Locations Nepal Flooding Midwest Flooding
Similar Disasters - Nearby Locations Santa Rosa Wildfire Woolsey Fire
Different Disasters - Distant Locations Hurricane Matthew Nepal Flooding
Different Disasters - Nearby Locations Hurricane Matthew Mexico Earthquake
Temporal Portugal Wildfire (Pre-Disaster) Portugal Wildfire
Semantic Shift
EuroSAT
Unseen Class
(Repeated for 10 Classes)
9 out of 10 classes (e.g., Annual Crop, Forest, Herbaceous Vegetation, Highway, Industrial, Pasture, Permanent Crop, Residential, River) 1 holdout class (e.g., Sea/Lake)
Table 2: Performance metrics comparing the oracle classifier $g_{oracle}$ and the surrogate classifier $g^{*}$ for EuroSAT and xBD datasets across various experimental setups. The $g_{oracle}$ classifier acts as an upper bound for the $g^{*}$, which is the proposed method. The † notation indicates that over 10 measurements, the difference was not found to be statistically significant ($p<0.05$). All conditions are post-event except where specified.
Type
Dataset
[…]
$g_{oracle}$ $g^{*}$ $g_{oracle}$ $g^{*}$
Covariate Shift
EuroSAT
Eastern Europe Western Europe $0.91\pm 0.04$ $0.89\pm 0.06$ $0.32\pm 0.11$ $0.37\pm 0.10$
xBD Midwest Flooding Nepal Flooding $1.00\pm 0.00$ $1.00\pm 0.00$ $0.01\pm 0.02$ $0.01\pm 0.02$
Woolsey Fire Santa Rosa Wildfire $0.94\pm 0.02$ $0.93\pm 0.02$ $0.23\pm 0.08$ $0.27\pm 0.09$
Nepal Flooding Hurricane Matthew $0.99\pm 0.01$ $0.98\pm 0.01$ $0.06\pm 0.04$ $0.07\pm 0.02$
Mexico Earthquake Hurricane Matthew $0.96\pm 0.02$ $0.94\pm 0.03$ $0.11\pm 0.05$ $0.19\pm 0.08$
Portugal Wildfire (Pre) Portugal Wildfire $0.99\pm 0.00$ $0.95\pm 0.01$ $0.06\pm 0.04$ $0.23\pm 0.20$
Santa Rosa Wildfire Woolsey Fire $0.94\pm 0.02$ $0.93\pm 0.02$ $0.23\pm 0.08$ $0.27\pm 0.09$
Semantic Shift EuroSAT Forest Remaining 9 classes $0.99\pm 0.01$ $0.97\pm 0.02$ $0.08\pm 0.04$ $0.08\pm 0.05$
Herbaceous Vegetation $0.93\pm 0.03$ $0.88\pm 0.07$ $0.26\pm 0.11$ $0.29\pm 0.14$
Highway $0.63\pm 0.08$ $0.56\pm 0.06$ $0.78\pm 0.13$ $0.88\pm 0.10$
Industrial $0.97\pm 0.02$ $0.96\pm 0.02$ $0.13\pm 0.14$ $0.22\pm 0.07$
Pasture $0.98\pm 0.01$ $0.95\pm 0.01$ $0.09\pm 0.06$ $0.28\pm 0.19$
Permanent Crop $0.92\pm 0.03$ $0.89\pm 0.03$ $0.26\pm 0.08$ $0.33\pm 0.11$
Residential $0.87\pm 0.03$ $0.82\pm 0.06$ $0.46\pm 0.21$ $0.43\pm 0.15$
River $0.87\pm 0.04$ $0.81\pm 0.06$ $0.44\pm 0.14$ $0.57\pm 0.10$
Sea-Lake $1.00\pm 0.00$ $1.00\pm 0.00$ $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00
Annual Crop $0.92\pm 0.03$ $0.91\pm 0.04$ $0.25\pm 0.10$ $0.32\pm 0.16$
5 Experimental Setup
In this section, we describe the experimental setups designed to evaluate TARDIS under controlled distribution shifts, which mimic real-world geospatial deployment challenges. Specifically, we use the EuroSAT [9] dataset for patch-level classification and the xBD [7] dataset for semantic segmentation. EuroSAT contains 27,000 labeled images across ten land use and land cover classes: Annual Crop, Forest, Herbaceous Vegetation, Highway, Industrial, Pasture, Permanent Crop, Residential, River, and Sea/Lake. The xBD dataset includes over 1 million labeled samples for building damage assessment from satellite imagery, covering multiple disaster types such as hurricanes, floods, and wildfires. For our experiments, we formulate xBD as a building footprint segmentation task with two classes (building footprint and background). EuroSAT and xBD provide a suitable test-bed for our OOD method due to their differences in volume, task, sensor type, geospatial distribution, and temporal and spatial characteristics, allowing a thorough evaluation under conditions commonly encountered in geospatial deployments.
To introduce controlled distribution shifts, we reformulate the train, validation, and test splits of EuroSAT and xBD so that the model, when trained on the modified training set and evaluated on the adjusted test set, encounters shifts similar to those seen in geospatial model deployments. We categorize these shifts as covariate shift and semantic shift. Covariate shift occurs when the input data distribution changes between training and testing. This can introduce selection bias, as covariates in the training set may still appear in the test set to some extent, potentially skewing robustness evaluation. To address this, we also introduce semantic shift experiments where a class is held out during testing, creating an unseen semantic class. For our experiments, covariate shifts include spatial variance (varying spatial locations), temporal variance (varying time of data collection), and environmental variance (varying landcover or disaster types). For semantic shifts in the EuroSAT dataset, we train the model on nine out of ten classes and test on the holdout class, repeating this process so that each class serves as the holdout class once. Details of the experimental setups are provided in Table 1.
The setups described above assume clear definitions of ID (training set) and OOD (test set), which presupposes access to test-time distribution shifts, an assumption that is not feasible in real-world deployments. To evaluate our method’s performance under unknown test-time distribution shifts, we fit and evaluate the binary classifier $g$ in two distinct ways. In the first approach, we bypass surrogate label assignment (Step 3 in Figure 2), using the well-defined ID and OOD samples to fit and evaluate the classifier, referred to as $g_{\text{oracle}}$. In the second approach, we treat the OOD labels in the test set as unknown, simulating real-world conditions where these labels are unavailable. This method uses surrogate label assignment to reassign unknown labels into surrogate ID and surrogate OOD. Parameter estimation and evaluation on these reassigned labels yield a classifier denoted as $g^{*}$, representing our proposed method. The underlying assumption is that $g_{\text{oracle}}$ provides an upper bound on performance for $g^{*}$, as $g_{\text{oracle}}$ is trained with clearly separated ID and OOD samples, whereas $g^{*}$ is trained on surrogate ID and surrogate OOD samples that are generated through our surrogate label assignment process.
Figure 3: Geographical distribution of ID and WILD sets, containing 500 and 1200 samples, respectively. The ID set is sampled from the FTW dataset training set, while the WILD set is randomly sampled from the Microsoft Planetary Computer. Each Sentinel-2 patch, provided in two different time frames (planting and harvesting). The model $f$ takes both Sentinel-2 images from different seasons as input, and the predictions are shown on the right. The bar indicates the probability output of $g$ for each input pair, with a higher probability (closer to 1) suggesting a higher likelihood of the pair being OOD. On the map, probability values are thresholded at 0.5.
6 Experimental Results
In this section, we present the ablation studies and discuss the results of the surrogate sample assignment benchmark, where we evaluate the performance of $g^{*}$. Note that ID task performance is not reported, as it remains unaffected by our method.
Ablation Studies. We conduct benchmark studies on the $g_{\text{oracle}}$ classifier to explore the factors influencing TARDIS’s ability to detect test-time distribution shifts. This ability depends significantly on the layer from which internal activations are extracted and the downsampling method applied. To address this, we test different layers and compare their performance. For downsampling, we experiment with several methods, including mean and standard deviations, average pooling, max pooling, and PCA-based reduction. The max pooling-based downsampling method achieves the highest performance, likely due to its ability to retain the most salient activation patterns, which we argue is important for effective OOD detection. After selecting the optimal layer and downsampling method, we evaluate various classifiers: KNeighbors, GaussianNB, DecisionTree, ExtraTrees, LogisticRegression, SVC, RandomForestUnbalanced, RandomForest, AdaBoost, and GradientBoosting. Results indicate that Logistic Regression provides the best tradeoff between classification performance and wall time. We then compare $g^{*}$ to $g_{\text{oracle}}$ as we tune the parameters $K$ and $T$ required for surrogate label assignment. To determine $K$ and $T$, we use a Tree-structured Parzen Estimator for sampling. Specifically, we set the search boundaries for $K$ (the number of clusters) between 2 and 0.3 $\times$ M, where $M$ is the total number of samples, and for $T$ (the fraction of in-distribution samples within a cluster, below which the cluster is labeled as surrogate OOD) between 0.01 and 0.2. These values were empirically determined based on preliminary tuning for optimal performance. We then run 20 independent experiments and select the best-performing $(K,T)$ pair. Fixing $T$ based on the hyperparameter search results, we observe a recurring pattern that enables us to fix $K$ to 0.3 $\times$ M across all setups. This choice is driven by the observation that the classifier $g^{*}$ reaches performance levels close to $g_{\text{oracle}}$ when $K$ is set to 30% of the total training samples. This trend holds across both datasets, with the xBD disaster scenario (semantic shift) and EuroSAT Pasture experiment (covariate shift) showing similar behavior. In both cases, classifier performance improves as more clusters are incorporated, approaching the performance of $g_{\text{oracle}}$. We present the results of TARDIS across these controlled distribution shifts in the Supplementary Material.
Surrogate Sample Assignment Benchmark. The quantitative results for $g^{*}$ and $g_{\text{oracle}}$ across 16 experimental setups are shown in Table 2. Across all setups, $g^{*}$ closely matches the performance of $g_{\text{oracle}}$, which represents the upper bound for our method. Specifically, in 13 setups for AUROC and 17 setups for FPR95, the difference between $g^{*}$ and $g_{\text{oracle}}$ is not statistically significant. In the remaining setups, while a small performance difference exists, the proposed method largely recovers the theoretical upper bound. This minimal, and often statistically insignificant, performance gap demonstrates that TARDIS can assign surrogate ID and OOD labels with high performance, even without access to test-time distribution during training.
A key takeaway from these experiments is the method’s ability to disentangle WILD samples with unknown test-time distributions into surrogate ID and OOD samples. Notably, under unknown distribution settings, it achieves performance comparable to cases with known ID and OOD labels, highlighting its strength in reliable OOD detection without prior distribution knowledge, making it suitable for real-world geospatial applications.
7 OOD Detection Goes Global: Real-World Deployment
We use the Fields of the World (FTW) pre-trained models [11] to demonstrate the capabilities of TARDIS in a large-scale deployment scenario. FTW is a geographically diverse dataset designed for agricultural field segmentation, covering 24 regions across Europe, Africa, Asia, and South America: Austria, Belgium, Brazil, Cambodia, Corsica, Croatia, Denmark, Estonia, Finland, France, Germany, India, Kenya, Latvia, Lithuania, Luxembourg, Netherlands, Portugal, Rwanda, Slovakia, Slovenia, South Africa, Spain, Sweden, and Vietnam. The dataset contains approximately 70,000 samples, each consisting of multi-date, multi-spectral Sentinel-2 satellite patches paired with three-class semantic segmentation masks (field, field boundary, and background). The task involves segmenting these classes using a pair of Sentinel-2 images — one for the planting season and one for the harvesting season — as input. Field boundary data is crucial for global agricultural monitoring, however training large scale models to segment field boundaries from satellite imagery presents significant challenges due to the geographic diversity of fields, varying crop cycles, and agro-climatic conditions, all of which which introduce substantial distribution shifts. This complexity, along with corresponding distribution shifts found in real world imagery, makes the FTW dataset ideal for testing TARDIS’s ability to diagnose model performance during inference. Additionally, the multi-date nature of the dataset is particularly suitable for evaluating segmentation models that must handle spatiotemporal variations in satellite imagery.
Sampling ID Set. To form the ID set, we sample 50 patches from each country represented in the training set, ensuring a geographically diverse ID dataset that closely matches the data on which the model was originally trained.
Sampling WILD Set. We collect multispectral Sentinel-2 satellite images using the Microsoft Planetary Computer [22], which is freely accessible. The images are processed to Level-2A (bottom-of-atmosphere) and stored in cloud-optimized GeoTIFF (COG) format. We use four spectral bands: Red (B04), Green (B03), Blue (B02), and Near-Infrared (B08), each with a spatial resolution of 10 meters per pixel. We randomly query 1200 Sentinel-2 scenes from those over land and with cloud coverage $<10\%$. Then, from each scene, we generate random patches of shape 256 $\times$ 256 $\times$ 4 pixels (filtering out samples with $\geq 10\%$ zero or NaN pixels) and query both the planting and harvesting seasons for the same locations. The planting and harvesting dates are fixed based on the hemisphere: April 1 to June 30 and September 1 to November 30 for the Northern Hemisphere, and October 1 to December 31 and March 1 to May 31 for the Southern Hemisphere. The two patches are concatenated along the channel dimension, resulting in a patch of shape 256 $\times$ 256 $\times$ 8 pixels. Following the pre-trained model’s convention, the concatenated patch is upsampled by a factor of 2 to 512 $\times$ 512 $\times$ 8, processed by the model, and then downsampled back to the original dimensions. For surrogate label assignment, we fix the hyperparameters $K$ and $T$ to $0.3\times 1200$ and 0.1, respectively, as suggested by the benchmark study.
The geographical distribution of ID and WILD samples is illustrated in Figure 3, where WILD samples are further classified into ID and OOD categories. A notable pattern is that samples located in arid biomes — such as the deserts in Inner Australia, the Sahara, and Patagonia — and polar regions, including Icelandic glaciers and the South Pole, are more likely to be assigned as OOD. This can be attributed to the ID samples predominantly representing mesic environments, which are areas with moderate moisture levels and managed landscapes suitable for agriculture. In contrast, arid biomes and polar regions exhibit extreme environmental conditions that differ markedly from the mesic environments of the ID samples. This ecological dissimilarity results in these biomes being classified as surrogate OOD, highlighting the model’s sensitivity to environmental context and distribution shifts.
To evaluate the scalability of our method, we measure the time required to classify the internal activation of a WILD sample as surrogate ID or OOD using the classifier $g$, implemented as a logistic regressor. An input patch of shape $256\times 256$ pixels, with a spatial resolution of 10 meters per pixel, covers an area of $6.5536\,\text{km}^{2}$ and takes 0.003 seconds to process. Applied to the entirety of mainland Africa, which spans $29.77\times 10^{6}\,\text{km}^{2}$, the method could be deployed on a single CPU in approximately 15 days. In a distributed GPU setup, this process could be completed within hours (a fraction of the time required for $f$ model inference over a similar area). This is crucial as such a model would allow practitioners to see, spatially, where the field boundary model itself is performing in distribution versus not (and where the corresponding outputs are likely irrelevant). These results demonstrate the practicality of the proposed method for deployment in low-data regimes, where geospatial models are typically less robust, and highlight its potential for improving model understanding in these challenging settings.
We further evaluate the reliability of the OOD classifier $g$ on the training set by comparing the performance of the $f$ model on the FTW dataset’s test set with the probability scores from $g$. We observe a correlation between the performance of the $f$ model on the test set and low skewness in the output of $g$. This indicates that when the $f$ model performs poorly, the OOD classifier $g$ reflects this by producing unimodal OOD scores, as shown by low skewness values (see Figure 4) for countries such as Portugal, Cambodia, and Vietnam.
Figure 4: Relationship between model performance and OOD score skewness. The figure illustrates the correlation between low performance of the $f$ model on the FTW test set and low skewness in the OOD classifier $g$’s score distribution. When the $f$ model performs poorly, the OOD classifier $g$ produces more OOD scores, as reflected by lower skewness values.
8 Conclusion
We present TARDIS, which effectively assigns surrogate ID or OOD labels to samples with unknown distributions by clustering the representation space. These surrogate labels are used to train a binary classifier for OOD detection. We do not aim to optimize OOD detection performance but instead propose a method designed for real-world applications where the distribution is unknown, maintaining ID task performance is critical, and computationally expensive methods are impractical. It achieves performance close to the theoretical upper limit in assigning surrogate ID and OOD samples in 13 out of 17 experiments. We demonstrate TARDIS can be deployed at scale, with the potential to enhance the robustness and trustworthiness of geospatial models as well as provide insights into distribution shifts in global applications.
Supplementary Material
In this supplement, we first detail the datasets and models used, followed by a discussion of the introduced distribution shifts and their design rationale. Next, we evaluate the impact of various design choices, including layer selection, downsampling methods, classifiers, and surrogate label assignment hyperparameters. Finally, we present additional experimental results with detailed visualizations and performance metrics to provide deeper insights into the behavior and performance of our method.
9 Datasets and Model Details
9.1 EuroSAT
EuroSAT [9] is a scene classification dataset derived from Sentinel-2 satellite images, covering various locations across Europe. It contains 27,000 images labeled into ten land-use and land-cover classes: Annual Crop, Forest, Herbaceous Vegetation, Highway, Industrial, Pasture, Permanent Crop, Residential, River, and Sea/Lake. The images have a spatial resolution of 10 meters.
We use a ResNet50 model pre-trained on ImageNet, modified to accept 13 input channels corresponding to Sentinel-2 spectral bands. The model is fine-tuned with a learning rate of 0.0001 and a batch size of 128. Training runs for up to 100 epochs with early stopping after 5 epochs of no improvement. Input images are normalized using channel-wise mean and standard deviation statistics.
9.2 xBD
xBD [7] is a semantic segmentation dataset for building damage assessment from satellite imagery. The dataset, collected from Maxar’s Open Data Program, has images with a spatial resolution below 0.8 meters. It includes pre- and post-disaster images of hurricanes, floods, wildfires, and earthquakes, making it suitable for evaluating temporal and semantic shifts.
We simplify the damage assessment task into binary segmentation by reassigning damage levels: background (0) and levels 1-2 are grouped, while levels 3-4 form a high-damage class. This minimizes concept drift and ensures a balanced evaluation. We train a U-Net model with a ResNet50 backbone, pre-trained on ImageNet and configured for 3 input channels. Training uses a batch size of 32, a learning rate of 0.001, and runs for up to 50 epochs with early stopping after 5 epochs of no improvement. We reserve 10% of the data for validation and normalize the input images by dividing pixel values by 255.
9.3 FTW
We follow the practices of the original study and use a U-Net model with an EfficientNet-B3 backbone for semantic segmentation on the FTW dataset. The model is configured with 8 input channels and outputs 3 classes: background, field, and field-boundary. We use class weights of [0.04, 0.08, 0.88] to address class imbalance. The learning rate is set to 0.001, and the loss function is cross-entropy. The number of filters is set to 64, and neither the backbone nor the decoder is frozen during training. We set the patience for early stopping to 100 epochs. The images are normalized by dividing pixel values by 3000.
For the classifier, we use logistic regression with a maximum number of iterations set to 500. We train the classifier with 500 ID samples and 1200 WILD samples. The number of clusters is set to 150, calculated as 0.3 times the total number of WILD samples. To reassign labels, we use an ID fraction threshold of 0.1, meaning that a cluster is assigned as OOD if ID samples comprise less than 10% of the total samples in the cluster. The values of 0.3 and 0.1 are determined based on empirical observations gathered from extensive experiments on the xBD and EuroSAT datasets.
Figure 7 provides a visual illustration of the input samples from the WILD set, where the distribution is unknown. It displays the input Sentinel-2 image pair (Window A and Window B) alongside the OOD classifier $g$’s prediction scores and the DL model $f$’s predictions.
9.4 Introducing Distribution Shifts to EuroSAT and xBD
The combination of EuroSAT and xBD provides a diverse testbed for evaluating distribution shifts. EuroSAT represents regional imagery at medium spatial resolution, while xBD provides global imagery at very high resolution. Their differences in acquisition times, sensor parameters, processing levels, and the tasks they cover—land-cover classification (EuroSAT) and building detection (xBD)—make them complementary. Additionally, EuroSAT focuses on patch-level classification, while xBD involves pixel-level segmentation, enabling evaluations across different problem dimensions.
To evaluate our method, we introduce two types of distribution shifts: covariate and semantic (described in Table  1). Our approach assumes that purposefully rearranging dataset splits creates measurable shifts between training and testing sets, driven by the logic of the split design.
EuroSAT Distribution Shifts.
Figure  6 shows one example from each of EuroSAT’s 10 classes, which differ spatially and semantically. For covariate shifts, we split the dataset by longitude at the midpoint of its spatial extent, using the western half for training and the eastern half for testing. This creates a shift based on spatial proximity.
For semantic shifts, we train the model on 9 classes and test it on the hold-out class, repeating this process for all classes. This ensures the model faces unseen scenarios during testing, providing a robust evaluation of its ability to handle semantic shifts.
xBD Distribution Shifts.
Figure  5 illustrates the pre- and post-disaster image pairs in the xBD dataset. Temporal shifts arise from changes occurring between pre- and post-disaster images, while spatial and thematic shifts reflect differences in how disasters impact regions and leave varying degrees of visible marks. Using these inherent characteristics, we design covariate shift experiments for xBD.
10 Design Choices
To better understand the impact of various design choices on the performance of our OOD detection method, we conduct a series of ablation studies. Specifically, we explore four key factors: (1) the choice of layer from which to extract feature representations (Section 10.1), (2) the method used to downsample these feature maps (Section 10.2), (3) the type of binary classifier $g$ used to distinguish between surrogate-ID and surrogate-OOD samples (Section 10.3), and (4) the selection of hyperparameters k and T for surrogate label assignment (Section 10.4).
10.1 Which Layer?
Selecting the appropriate layer for activation extraction is crucial for accurate OOD detection. Prior works have emphasized the importance of this choice. For example, ASH achieves optimal performance on later layers like the penultimate layer, as earlier layers suffer from significant performance degradation during pruning [4]. Similarly, ReAct performs best on the penultimate layer, where more distinctive patterns between ID and OOD data emerge [24]. NAP-based OOD detection further highlights the variability in layer effectiveness, dynamically selecting top-performing layers based on validation accuracy [20]. Consistent with these findings, we observe that no single layer is universally optimal across all settings.
We benchmark FPR95 scores for OOD detection across the first convolutional layer, eight randomly selected intermediate layers, and the last convolutional layer. As shown in Table 3 for the EuroSAT dataset and Table 4 for the xBD dataset, layer performance varies significantly. While late layers often perform well, early and middle layers frequently give competitive results, depending on the dataset and task. Based on these findings, we select the best-performing layer for each experiment.
For the large-scale FTW dataset, the lack of distribution shift information prevents evaluation of layer-specific performance for OOD detection. Therefore, based on the observation that many layer benchmarks perform optimally for middle layers, we select a middle convolutional layer, specifically ‘decoder.blocks.0.conv1‘ from the U-Net model with an EfficientNet-B3 backbone.
Figure 5: Examples from the xBD dataset, illustrating pre- and post-disaster images. These samples demonstrate the temporal and semantic differences between pre- and post-disaster scenes, highlighting the challenges posed by distribution shifts.
Figure 6: Examples from the dataset, with one sample from each class. These images highlight the spatial and semantic distinctions across classes.
Figure 7: The input samples are from the WILD set, where the distribution is unknown. The figure shows Sentinel-2 images at two different times (planting season and harvesting season — Window A and Window B). When these windows are fed together into $f$, the model outputs both the segmentation prediction and the OOD classifier $g$’s prediction score.
10.2 Which Downsampling Method?
Having identified the layer to extract internal activations from, the next step is to look into the effect of downsampling these activations, which can reduce computational complexity and noise while retaining essential features for OOD detection. We explored four methods:
  1. 1.
    Mean and standard deviation (Mean Std): Computes the mean and standard deviation across the spatial dimensions (H, W) for each channel, providing two descriptive statistics per feature channel.
  2. 2.
    Average pooling (Avg Pool): Global average pooling was applied, reducing the activation to a single representative value per channel by averaging all spatial values.
  3. 3.
    Max pooling (Max Pool): Uses global max pooling to retain the maximum value from each spatial dimension, capturing the most prominent feature in each channel.
  4. 4.
    PCA-based reduction (PCA): Applies Principal Component Analysis (PCA) to reshape the activation map into a vector and projects it into a lower-dimensional space with 10 components.
We summarize the OOD detection performance across all experiments on the EuroSAT and xBD datasets under different downsampling methods in Table 5, using the FPR95 metric. Max pooling consistently achieves the best performance across the majority of experiments, making it the preferred approach. We attribute its performance to its ability to retain the most prominent features in each channel, filtering out less significant information. This focus on salient patterns likely enhances the OOD classifier’s capacity to distinguish between ID and OOD samples.
10.3 Which Classifier?
The next key design choice is the selection of the binary classifier $g$, used to distinguish between surrogate-ID and surrogate-OOD samples based on their feature representations. The results, summarized in Table 6, report the mean performance across all experimental measurements along with the standard error of the mean to represent confidence intervals. We select Logistic Regression as it provides the best tradeoff between classification accuracy and prediction time. This balance is essential for scaling up the method, where both efficiency and accuracy are critical.
10.4 Surrogate Label Assignment: Hyperparameter Search for k and T
TARDIS relies on a clustering-based approach in the activation space to assign surrogate ID and surrogate OOD labels. This process requires selecting two key parameters: the number of clusters (k) to segment the activation space, and the ID fraction threshold (T), which determines whether a cluster is assigned as surrogate ID or surrogate OOD. Clusters with an ID fraction above T are assigned as surrogate ID, and those below T are assigned as surrogate OOD.
The underlying assumption is that samples with similar distributions lie closer in the activation space than those from dissimilar distributions. Effectively segmenting the activation space is critical, as the distributions of WILD samples are unknown during deployment. Segmenting nearby activations and correctly assigning clusters to surrogate ID or OOD makes the choice of k and T crucial for accurate OOD detection.
To develop insights into selecting k and T, we conduct controlled experiments on EuroSAT and xBD, where ID and OOD labels are known. In these experiments, we treat OOD labels as WILD and apply our clustering-based surrogate label assignment logic. By holding back the ground-truth WILD labels, we simulate real-world conditions while being able to evaluate the results against known labels.
The primary goal is to understand how to choose k and T, and whether there are patterns we can extrapolate to real-life deployment. For this, we calculate the ratio of k to the total number of training samples and evaluate its effect on OOD detection performance (Accuracy, FPR95, and AUROC). We plot these metrics against the ratio of k/total training samples, increasing k until the ratio reaches 1. Theoretically, OOD detection improves with more clusters as this enables finer-grained segmentation of the activation space, reducing the risk of including anomalies in ID clusters.
To establish a theoretical maximum (oracle performance), we also evaluate OOD classification with known ID and OOD labels, bypassing the need for clustering. This oracle performance is represented by horizontal dashed lines in Figures 8 and 9 (upper plots). The results for two representative experiments—one from EuroSAT and one from xBD—since all experiments show similar trends. We observe that the performance approaches the oracle boundaries when k is approximately 0.3 times the total number of training samples. While performance improves as k increases, a trade-off is required between performance and walltime as well as computational complexity. Based on this trade-off, we set k to 0.3 for all experiments, including the large-scale deployment on FTW. Furthermore, we observe that our method is not highly sensitive to T. As a result, we fix T to 0.1 for all experiments, which is the value used in this initial investigation. We use the Optuna library to implement a Bayesian-based search algorithm. The composite objective function, which we minimize to determine the optimal number of clusters and ID fraction threshold, is detailed in Section 4.
Lastly, the gradual improvement in OOD detection performance with increasing k supports our assumption that samples with similar distributions lie closer in the activation space than those with dissimilar distributions. The absence of degradation in performance further underscores the importance of activation-level clustering as a reliable proxy for domain estimation based on neighboring samples.
We set k and T as described and use t-SNE in the lower plots of Figures 8 and 9 to reduce the dimensionality of the activation spaces to 2D for visualization. When ID and OOD labels are known, the t-SNE plots show that only a small fraction of labels changes from the original labels. This demonstrates the effectiveness of the surrogate label assignment process described above.
11 Additional Experimental Results
In Figure 10, we show the predictions of the DL model $f$ and the OOD classifier $g$, along with the ground truth class and distribution annotations for the EuroSAT experiment, where Forest serves as the OOD class. The model $f$ trains on 9 classes (excluding Forest) and tests on Forest. The first row shows correct predictions by $f$, while the second row shows incorrect predictions. Even when $f$ makes misclassifications, $g$ accurately quantifies the distribution shifts in most cases. The performance of $f$ on the test set is not directly measurable since the test uses a single unseen class. We report the performance of $g$ as: Accuracy = 0.9325, ROC AUC = 0.9886, and FPR95 = 0.0619.
For xBD, we present results where $f$ trains on Hurricane Matthew (ID, Figure 11) and tests on Mexico Earthquake (OOD, Figure 12). Comparing the input images and masks between ID and OOD shows that even when $f$ performs suboptimally, $g$ effectively quantifies the distribution shifts. The performance of $f$ on the test set is: Multiclass Accuracy = 0.7690 and Multiclass Jaccard Index = 0.6248. We attribute $f$’s suboptimal prediction performance to the significant distribution shift between the training (Hurricane Matthew) and testing (Mexico Earthquake) datasets. The performance of $g$ is: Accuracy = 0.9806, ROC AUC = 0.9986, and FPR95 = 0.0.
Experiment 2/217 8/217 16/217 38/217 43/217 48/217 118/217 139/217 199/217 211/217
Forest 0.0625 0.01 0.00 0.00 0.00 0.00 0.0078 0.0156 0.0391 0.0391
HerbVeg 0.2857 0.22 0.3095 0.22 0.2778 0.1429 0.07 0.1032 0.2460 0.2778
Highway 0.8319 0.5462 0.6218 0.3529 0.3613 0.21 0.12 0.0840 0.1765 0.2437
Industrial 0.2406 0.01 0.0376 0.0226 0.0226 0.00 0.0150 0.0226 0.0376 0.0075
Pasture 0.1288 0.0909 0.12 0.1364 0.0985 0.03 0.0833 0.0227 0.1212 0.2273
PermCrop 0.3554 0.2975 0.3140 0.2397 0.2314 0.14 0.12 0.1322 0.2066 0.1653
Residential 0.2960 0.00 0.0160 0.0240 0.00 0.00 0.00 0.0160 0.0400 0.0480
River 0.4688 0.07 0.2031 0.03 0.0938 0.0234 0.0078 0.0078 0.0625 0.0859
SeaLake 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
AnnualCrop 0.2879 0.00 0.0303 0.0606 0.0530 0.02 0.0379 0.0152 0.0682 0.0758
SpatialSplit 0.3182 0.20 0.4773 0.15 0.1970 0.0909 0.2197 0.2273 0.6364 0.7652
Avg ± Stdev 0.2978 ± 0.2223 0.1313 ± 0.1729 0.1936 ± 0.2133 0.1124 ± 0.1176 0.1214 ± 0.1264 0.0597 ± 0.0741 0.0620 ± 0.0696 0.0588 ± 0.0712 0.1486 ± 0.1801 0.1760 ± 0.2185
Table 3: FPR95 scores for OOD detection for experiments on EuroSAT dataset across the first convolutional layer, eight randomly selected layers, and the last convolutional layer. Notation in the header (e.g., X/Y) refers to the ’layer number / total number of layers.’ The last row, labeled ’Avg ± Stdev,’ provides the mean ± standard deviation of the scores for each layer across all experiments.
Experiment 3/223 60/223 112/223 146/223 148/223 162/223 176/223 187/223 208/223 216/223
Nepal Flooding
- Midwest Flooding
1.0000 0.09 0.6986 0.7534 0.6986 0.7397 0.46 0.8767 0.8356 0.8493
Santa Rosa Wildfire
- Woolsey Fire
0.5000 0.36 0.36 0.7222 0.6389 0.6389 0.6944 0.6111 0.58 0.9167
Hurricane Matthew
- Nepal Flooding
0.3023 0.06 0.3488 0.4419 0.4884 0.32 0.1163 0.3488 0.2791 0.4419
Hurricane Matthew
- Mexico Earthquake
0.11 0.1471 0.2353 0.6471 0.4118 0.50 0.3235 0.4706 0.6471 0.9706
Portugal Wildfire
(Pre-Post)
0.38 0.9583 0.3889 0.8472 0.8750 0.9167 0.9722 0.77 0.9861 0.8472
Mean ± Stdev 0.4585 ± 0.3343 0.3231 ± 0.3740 0.4063 ± 0.1735 0.6824 ± 0.1524 0.6225 ± 0.1818 0.6231 ± 0.2275 0.5133 ± 0.3316 0.6154 ± 0.2146 0.6656 ± 0.2686 0.8051 ± 0.2095
Table 4: FPR95 scores for OOD detection experiments on the xBD dataset across the first convolutional layer, eight randomly selected layers, and the last convolutional layer. Notation in the header (e.g., X/Y) refers to the ’layer number / total number of layers.’ The last row, labeled ’Avg ± Stdev,’ provides the mean and standard deviation of the scores for each layer across all experiments.
Experiment Avg Pool Mean Std Max Pool PCA
Forest 0.0859 0.4297 0.0234 0.8750
HerbaceousVegetation 0.2937 0.8651 0.2698 0.9921
Highway 0.7899 0.7311 0.7059 0.9412
Industrial 0.1880 0.2857 0.0526 0.9925
Pasture 0.0909 0.6970 0.2576 0.9924
PermanentCrop 0.3058 0.9008 0.4215 0.9669
Residential 0.2640 0.6640 0.2160 0.8480
River 0.4922 0.6172 0.1563 0.9922
SeaLake 0.0000 0.0313 0.0000 0.9766
AnnualCrop 0.2500 0.7576 0.0455 0.9924
SpatialSplit 0.3182 0.5379 0.3030 0.9848
Nepal Flooding - Midwest Flooding 0.0000 0.6575 0.9452 0.9726
Hurricane Matthew - Nepal Flooding 0.0233 0.9070 0.5581 0.9535
Hurricane Matthew - Mexico Earthquake 0.0588 0.9118 0.5882 1.0000
Portugal Wildfire Pre - Portugal Wildfire Post 0.9028 0.9861 0.8472 1.0000
Table 5: FPR95 scores for OOD detection across different downsampling methods. The table compares performances of average pooling, mean and standard deviation pooling, max pooling, and PCA for various experiments. Bold values indicate the best performance for each experiment, while italicized values represent the second-best performance.
[…]
KNeighbors 92.23 ± 0.81 86.85 ± 1.07 38.79 ± 2.02 73.00 ± 8.00
GaussianNB 84.28 ± 1.04 89.37 ± 0.91 32.89 ± 1.89 4.00 ± 1.00
DecisionTree 91.21 ± 0.93 77.43 ± 1.20 78.81 ± 2.41 2.00 ± 1.00
ExtraTrees 93.30 ± 0.67 91.29 ± 0.83 28.53 ± 1.94 12.00 ± 3.00
LogisticRegression 87.67 ± 1.00 93.33 ± 0.87 27.20 ± 1.98 3.00 ± 1.00
SVC 91.54 ± 0.90 94.26 ± 0.72 19.87 ± 1.85 67.00 ± 12.00
RandomForestUnbalanced 92.54 ± 0.82 91.24 ± 0.80 30.98 ± 1.95 9.00 ± 2.00
RandomForest 92.76 ± 0.75 91.11 ± 0.84 30.24 ± 1.91 8.00 ± 2.00
AdaBoost 92.85 ± 0.79 92.03 ± 0.82 29.84 ± 1.89 11.00 ± 3.00
GradientBoosting 92.92 ± 0.81 93.11 ± 0.85 30.47 ± 1.88 7.00 ± 2.00
Table 6: Benchmark results of classifiers $g$, including Accuracy, ROC_AUC, FPR95, and prediction time. Values are reported as mean ± SEM over all experiments on EuroSAT and xBD. Bold indicates the best performance, and italics indicate the second-best performance. Prediction time is reported in milliseconds (ms/sample).
Figure 8: EuroSAT Pasture experiment on surrogate label assignment. The upper plot shows the performance metrics (Accuracy, FPR95, AUROC) for the oracle classifier $g_{\text{oracle}}$ and the surrogate classifier $g^{*}$ as the ratio of clusters to training samples $k/\text{len}(X_{\text{train}})$ increases. As $k$ grows, $g^{*}$ gradually improves and approaches the performance of $g_{\text{oracle}}$. The lower plot visualizes the feature space before and after clustering, showing how original ID and OOD labels are reassigned to surrogate ID and OOD labels based on the clustering logic.
Figure 9: xBD Nepal Flooding-Midwest Flooding disaster experiment on surrogate label assignment. The upper plot shows the performance metrics (Accuracy, FPR95, AUROC) for the oracle classifier $g_{\text{oracle}}$ and the surrogate classifier $g^{*}$ as the ratio of clusters to training samples $k/\text{len}(X_{\text{train}})$ increases. As $k$ grows, $g^{*}$ gradually improves and approaches the performance of $g_{\text{oracle}}$. The lower plot visualizes the feature space before and after clustering, showing how original ID and OOD labels are reassigned to surrogate ID and OOD labels based on the clustering logic.
Figure 10: EuroSAT experiment with Forest as the OOD class. The figure shows predictions of the DL model $f$ and the OOD classifier $g$, along with the ground truth class and distribution annotations. The first row represents samples where $f$ makes correct class predictions, while the second row represents samples where $f$ makes incorrect predictions. For each sample, we report both the ground truth distribution and the predicted distribution from $g$.
Figure 11: xBD experiment with Hurricane Matthew as the ID samples. The figure shows the annotations and predictions of the DL model $f$ and the OOD classifier $g$. For each sample, we present $f$’s predicted class and $g$’s predicted distribution, along with the ground truth annotations.
Figure 12: xBD experiment with Mexico Earthquake as the OOD samples. The figure shows the annotations and predictions of the DL model $f$ and the OOD classifier $g$. For each sample, we present $f$’s predicted class and $g$’s predicted distribution, along with the ground truth annotations.

### v2

@@
 Abstract
-Training robust deep learning models is critical in Earth Observation, where globally deployed models often face distribution shifts that degrade performance, especially in low-data regions.
-Out-of-distribution (OOD) detection addresses this challenge by identifying inputs that differ from in-distribution (ID) data.
-However, existing methods either assume access to OOD data or compromise primary task performance, making them unsuitable for real-world deployment.
-We propose TARDIS, a post-hoc OOD detection method for scalable geospatial deployments.
-The core novelty lies in generating surrogate labels by integrating information from ID data and unknown distributions, enabling OOD detection at scale.
-Our method takes a pre-trained model, ID data, and WILD samples, disentangling the latter into surrogate ID and surrogate OOD labels based on internal activations, and fits a binary classifier as an OOD detector.
-We validate TARDIS on EuroSAT and xBD datasets, across 17 experimental setups covering covariate and semantic shifts, showing that it performs close to the theoretical upper bound in assigning surrogate ID and OOD samples in 13 cases.
-To demonstrate scalability, we deploy TARDIS on the Fields of the World dataset, offering actionable insights into pre-trained model behavior for large-scale deployments.
-The code is publicly available at https://github.com/microsoft/geospatial-ood-detection
-Figure 1: Illustration of the proposed method for OOD detection.
-Given a pre-trained model, a dataset of known ID samples, and WILD samples (containing both ID and OOD data), the method assigns surrogate ID or OOD labels to each WILD sample and fits a binary classifier $g$ (top row).
-During deployment, the classifier $g$ uses internal activations to predict whether an unseen sample is ID or OOD (bottom row).
+Training robust deep learning models is crucial in Earth Observation, where globally deployed models often face distribution shifts that degrade performance, especially in low-data regions.
+Out-of-distribution (OOD) detection addresses this by identifying inputs that deviate from in-distribution (ID) data.
+However, existing methods either assume access to OOD data or compromise primary task performance, limiting real-world use.
+We introduce TARDIS, a post-hoc OOD detection method designed for scalable geospatial deployment.
+Our core innovation lies in generating surrogate distribution labels by leveraging ID data within the feature space.
+TARDIS takes a pre-trained model, ID data, and data from an unknown distribution (WILD), separates WILD into surrogate ID and OOD labels based on internal activations, and trains a binary classifier to detect distribution shifts.
+We validate on EuroSAT and xBD across 17 setups covering covariate and semantic shifts, showing near-upper-bound surrogate labeling performance in 13 cases and matching the performance of top post-hoc activation- and scoring-based methods.
+Finally, deploying TARDIS on Fields of the World reveals actionable insights into pre-trained model behavior at scale.
+The code is available at https://github.com/microsoft/geospatial-ood-detection
+Figure 1: Overview of the proposed OOD detection method.
+Given a pre-trained model, ID samples, and WILD samples (from unknown distributions), TARDIS assigns surrogate ID/OOD labels to WILD samples using the ID set and fits a binary classifier $g$ (top row).
+During deployment, $g$ uses internal activations of unseen samples to predict whether they are ID or OOD (bottom row).
 1 Introduction
-Deep learning (DL) models have demonstrated remarkable capabilities across various domains but often exhibit overconfidence in their predictions, even when confronted with data that diverges from their training distribution [25, 10, 19].
+Deep learning models have demonstrated remarkable capabilities across various domains but often exhibit overconfidence in their predictions, even when confronted with data that diverges from their training distribution [31, 13, 24].
 This overconfidence arises from the assumption that inference data will follow the same independent and identically distributed (i.i.d.) properties as the training data.
@@
 This degradation poses a critical challenge, especially in applications where incorrect predictions can have severe consequences.
-In Earth Observation (EO), specifically with satellite imagery, distribution shifts are even more common due to variability in sensor types, acquisition angles, temporal changes, spatial variations, and landscape alterations caused by events such as natural disasters.
-This variability necessitates unique, tailored approaches for satellite imagery [21].
-The nature of distribution shifts between training and inference data can be classified as near-distribution or far-distribution shifts, each posing unique challenges.
-For example, a model trained on data from a certain season and region may perform poorly when applied to other seasons in the same region or the same season in different regions [13], representing a far-distribution shift.
-Even models trained on data covering multiple seasons and regions may perform suboptimally under specific acquisition or atmospheric conditions, such as cloud cover, sensor faults, or artifacts, indicative of a near-distribution shift.
+In Earth Observation (EO), specifically with satellite imagery, distribution shifts between training and inference commonly occur in the form of covariate and semantic shifts.
+Covariate shift refers to changes in the input distribution while keeping the task or label space fixed, whereas semantic shift involves changes in the label distribution or the meaning of inputs.
+A specific case of covariate shift is known as subpopulation shift [17], where the overall data distribution changes due to different proportions of underlying subgroups, even though the subgroups themselves remain the same.
+In EO, this arises when certain seasons, regions, or acquisition conditions are underrepresented during training but appear more prominently at test time.
+The magnitude of distribution shifts can further be characterized along a spectrum from near-distribution to far-distribution shifts.
+Near-distribution shifts may occur within the same satellite source due to temporal, geographic, or environmental variation—such as seasonal changes, sensor degradation, or atmospheric effects like cloud cover [16].
+Far-distribution shifts involve more substantial differences, such as changes in satellite platforms, sensor types, or entirely different imagery domains, resulting in greater divergence from the training distribution.
+This variability necessitates tailored approaches for satellite-based machine learning models [26].
+Despite the vital role of robust models in EO, few studies explore OOD detection in this domain [6, 3, 2, 32, 7].
+These works often presume access to test-time distributions, adhere to closed-set assumptions, and/or degrade model performance on in-distribution tasks.
+This underscores the need for specialized OOD detection methods suited for large-scale EO deployment.
+A further challenge arises when applying geospatial models globally.
+Global geospatial models often suffer significant performance drops in low-data regions [1], such as sub-Saharan Africa [15].
+These issues emphasize the need for global models equipped to detect OOD samples during inference.
 Figure 2: The proposed framework consists of four key steps: (1) Sampling in-distribution (ID) and WILD samples; (2) Extracting internal activations from a pre-trained model $f$ for both ID and WILD samples; (3) Clustering the combined feature space and labeling WILD samples as surrogate-ID or surrogate-OOD; (4) Fitting a binary classifier $g$ on the labeled feature representations to distinguish between ID and OOD samples.
 The classifier $g$, during deployment, flags out-of-distribution inputs.
-Despite the critical importance of robust models in Earth Observation (EO), only a few studies address OOD detection in this domain [5, 3, 1].
-These studies typically assume access to test-time distributions, rely on closed-set assumptions, and often sacrifice model performance on in-distribution tasks.
-This highlights the need for tailored OOD detection methods suited for large-scale EO deployment.
-A further challenge arises when applying geospatial models globally.
-Many geospatial products, while measuring the same ground realities with high coverage, exhibit low levels of agreement [2].
-Additionally, global geospatial models often suffer significant performance drops in low-data regions, such as sub-Saharan Africa [12].
-These issues emphasize the need for global models equipped to detect OOD samples during inference.
-OOD detection capabilities in real-world applications enhance model reliability by identifying when input data deviates from the training distribution.
-They also support resource allocation and targeted data collection by pinpointing areas where the model may underperform, effectively guiding both computational and human resources.
-Additionally, OOD detection enables real-time adjustments, allowing models to remain reliable in dynamic conditions.
-This approach supports a human-in-the-loop strategy for critical decisions, fostering trust, transparency, and accountability.
-Overall, OOD detection is essential for managing large-scale EO deployments effectively.
-In this paper, we begin with an essential observation: OOD samples can trigger internal activation patterns that diverge from those of ID samples.
+In this paper, we build on well-established finding that OOD samples trigger internal activation patterns that diverge from ID samples [30, 25].
 Building on this insight and recognizing the challenges of deploying geospatial models at scale, we introduce the following key contributions:
-  • We propose TARDIS (Test-time Addressing of Distribution Shifts at Scale), a post-hoc OOD detection method that preserves the model’s performance on the main task and operates without requiring access to unknown test-time near-distribution shifts.
-  • We evaluate TARDIS on EuroSAT, a patch-classification dataset, and xBD, a semantic segmentation dataset, examining both covariate and semantic shifts across 16 experimental setups.
-For covariate shifts, we introduce geographical variance (spatial changes), temporal variance (seasonal changes), and environmental variance (landcover or disaster type changes).
-Semantic shifts are introduced by withholding a class during training and exclusively testing on that class, thus presenting new semantic categories.
-  • We demonstrate that TARDIS can be deployed at scale, providing actionable insights into the trustworthiness and robustness of geospatial models in real-world scenarios.
+  • We propose TARDIS (Test-time Addressing of Distribution Shifts at Scale), a lightweight post-hoc OOD detection method that preserves model performance and operates without access to labeled data from unknown test-time shifts.
+  • We evaluate TARDIS on EuroSAT (patch classification) and xBD (semantic segmentation) under near-distribution settings, analyzing covariate (geographical, temporal, environmental) and semantic (withheld class) shifts across 17 experimental setups.
+We also compare its performance against existing post-hoc OOD detection methods.
+  • We show that TARDIS scales effectively, providing actionable insights into the robustness and trustworthiness of geospatial models in real-world scenarios.
 2 Related Works
@@
 This reliance limits their effectiveness under evolving test-time conditions and shifting distributions.
+Another line of OOD detection research uses clustering in feature space to separate ID and OOD samples.
+These methods exploit structural patterns in learned representations rather than direct scoring or activation manipulation.
+Gulati et al.
+[8] improve over $k$-means by using non-negative kernel regression for soft clustering.
+Sinhamahapatra et al.
+[27] explore clustering in self-supervised latent spaces, showing its potential in low-label settings.
+Closely related in spirit, NG-Mix [5] proposes pseudo-anomaly generation for semi-supervised anomaly detection.
 The task of OOD detection is inherently challenging, and the EO domain is no exception due to the diverse and heterogeneous nature of satellite imagery.
 Despite the critical importance of OOD detection in EO, this area remains relatively underexplored, with only a limited number of studies addressing the issue.
-One such study develops a Dirichlet Prior Network to quantify distributional uncertainty in DL models for satellite image analysis [5].
+One such study develops a Dirichlet Prior Network to quantify distributional uncertainty in deep learning models for satellite image analysis [6].
 This method assumes Dirichlet distributions for ID samples and employs various setups where classes, color channels, and environmental features are alternately treated as ID or OOD.
 Another study frames anomaly detection as a cumulative open-set detection and location separation task [3].
-This approach also uses a Dirichlet prior, similar to the first study, to expand the representation space between ID (normal) and OOD (anomalous) samples by predicting the anticipated categorical distribution.
-The method is suitable for scenarios where pre-event images are either unavailable, subject to radiation differences, or not sufficiently recent to aid in detection.
-In this paper, we do not aim to optimize OOD detection performance but instead propose a method designed for real-world applications where the distribution is unknown, maintaining ID task performance is critical, and computationally expensive methods are impractical.
-Our focus is on handling near-distribution shifts, where data variations stem from the same satellite sources but exhibit changes over time, geography, or environmental factors.
-This contrasts with far-distribution shifts, where variations arise from entirely different data sources, such as different satellite models or natural images.
+This approach also uses a Dirichlet prior, like the first study, to expand the representation space between ID (normal) and OOD (anomalous) samples by predicting the anticipated categorical distribution.
+The method suits scenarios where pre-event images are unavailable, subject to radiation differences, or not recent enough to aid detection.
+In this paper, we do not aim to optimize OOD detection performance but propose a method suited for real-world applications where the distribution is unknown, maintaining ID task performance is critical, and expensive methods are impractical.
+Our method detects shifts within the same satellite sources.
+This type of near-distribution shift presents a greater challenge than far-distribution shifts, such as those between natural and satellite imagery, as the differences are more subtle and harder to disentangle.
+By tackling these nuanced shifts, our method makes OOD detection more practical for real-world geospatial applications.
 3 Problem Formulation
@@
 We define $g(\mathbf{z};\theta)=0$ if $\mathbf{x}\sim D_{\text{ID}}$, and $g(\mathbf{z};\theta)=1$ if $\mathbf{x}\sim D_{\text{OOD}}$.
-4 Method
-Fitting a classifier to distinguish between in-distribution (ID) and out-of-distribution (OOD) samples typically requires known ID and OOD labels, which is impractical in real-world deployments.
-This creates a need for surrogate ID and OOD labels for samples with unknown distributions, which we refer to as WILD samples.
-Our method, illustrated in Figure 1, operates with a given model, a set of known ID samples, and a set of WILD samples.
-It generates surrogate ID and OOD labels for the WILD samples, which are then used to train a binary classifier that can classify the distribution of incoming samples.
-First, we sample $M$ in-distribution samples from the ID set and $N$ WILD samples from the WILD set.
-Next, we pass both the ID and WILD samples through the frozen pre-trained model and extract activations from a specified layer.
-A pooling operation is applied to perform spatial downsampling, transforming the activations into one-dimensional vectors and combining the downsampled features.
-We then use k-means clustering to partition the combined features into $K$ clusters.
-For each cluster, we calculate the proportion of ID samples and assign surrogate labels based on a threshold $T$: if the proportion of ID samples in the cluster is greater than or equal to $T$, the surrogate label is 0; otherwise, it is 1.
-Here, $y_{\text{sur}}(\mathbf{z})$ denotes the surrogate label: 00 for surrogate ID and $1$ for surrogate OOD.
-With this labeled dataset, we fit a binary classifier $g(\mathbf{z};\theta)$, which maps a sample to a predicted label.
-During deployment, for a sample of unknown distribution, we perform activation extraction and classification as follows: $\hat{y}=g\left(\text{Downsample}\left(f(\mathbf{x})\right);\theta\right)$.
-The output can be interpreted as probability scores, allowing it to quantify the magnitude of domain shift or thresholded to produce binary values (00 for in-distribution and $1$ for out-of-distribution).
-Surrogate Label Assignment.
-The core novelty of TARDIS lies in its surrogate labeling generation step, which enables us to assign labels to WILD samples based on their proximity to ID samples in feature space.
-The assumption is that ID and OOD features are sufficiently distinct to allow effective separation through clustering.
-Specifically, we use k-means clustering to label WILD samples as surrogate OOD if WILD samples outnumber ID samples in a cluster, and as surrogate ID otherwise.
-Without access to true OOD data, ID samples serve as an indirect benchmark for distinguishing between ID and WILD samples.
-To optimize this process, we tune the number of clusters $K$ and threshold $T$ by balancing three objectives: entropy (to encourage homogeneity within clusters and distinct groups of surrogate labels), correct ID proportion (to maximize accurate identification of ID samples), and incorrect ID proportion (to minimize the misclassification of ID samples as OOD).
-The combined objective function minimizes average entropy and the misclassification rate of ID samples while maximizing correct ID classification, resulting in well-separated and accurately labeled groups.
-Figure 2 illustrates our method.
-First, ID samples are drawn from the model’s training set, while WILD samples, representing unknown distributions, are sourced from satellite images taken at any location and time (Step 1).
-Both ID and WILD samples are passed through a pre-trained model, and internal activations are extracted from a specified layer (Step 2).
-Using these activations, we assign surrogate labels to WILD samples, categorizing them as surrogate ID or surrogate OOD (Step 3).
-This surrogate labeling, central to our approach, decomposes WILD samples based on their distributional alignment with ID samples.
-Finally, these labeled feature representations are used to train a binary classifier that, during deployment, outputs probability scores or thresholded 0/1 predictions to classify samples as either ID or OOD (Step 4).
 Table 1: Overview of experimental setups for evaluating TARDIS under controlled distribution shifts.
@@
 9 out of 10 classes (e.g., Annual Crop, Forest, Herbaceous Vegetation, Highway, Industrial, Pasture, Permanent Crop, Residential, River) 1 holdout class (e.g., Sea/Lake)
+4 Method
+Training a classifier to distinguish in-distribution (ID) from out-of-distribution (OOD) samples typically requires labeled ID and OOD data, which is impractical in real-world settings.
+This motivates generating surrogate labels for samples with unknown distributions, referred to as WILD samples.
+Our method, illustrated in Figure 1, operates with a pre-trained model, known ID samples, and WILD samples.
+Using known ID samples, it assigns surrogate ID and OOD labels to WILD samples, which are then used to train a binary distribution classifier.
+TARDIS involves training a binary classifier to discriminate between ID and OOD samples based on internal activations, as shown in Figure 2.
+Given $M$ ID and $N$ WILD samples, we pass both through the frozen model and extract activations from a specified layer.
+After pooling for spatial downsampling, we obtain one-dimensional feature vectors, combine them, and assign corresponding ID or WILD labels.
+We shuffle this set and reserve 30% for validation.
+On the remaining 70%, we run $k$-means clustering to assign surrogate ID/OOD labels to WILD samples based on distance to ID samples.
+Using this labeled set, we train a binary classifier $g(\mathbf{z};\theta)$ to map activations to distribution labels.
+At deployment, we extract activations from a sample of unknown distribution and predict its label as $\hat{y}=g\left(\text{Downsample}\left(f(\mathbf{x})\right);\theta\right)$.
+The output can be interpreted as a probability of domain shift or thresholded for binary classification (00: ID, $1$: OOD).
+Surrogate Label Assignment.
+The core novelty of TARDIS lies in generating surrogate labels for WILD samples based on their proximity to known ID samples in feature space.
+We assume that ID and OOD features are separable enough for clustering to be effective.
+Specifically, we cluster the combined ID+WILD feature space into $k$ groups (Step 3 in Figure 2).
+For each cluster, if the proportion of ID samples is greater than or equal to the threshold $T$, all samples in that cluster receive label 0 (ID); otherwise, label 1 (OOD).
+Distance to ID samples thus serves as a proxy for assigning surrogate labels.
+Given $k$, $T$, and the feature vectors, we perform clustering and assign surrogate labels accordingly. To select the optimal parameters, we minimize the composite objective:
+$(k^{*},T^{*})=\arg\min_{k,T}\quad H(S)+P_{\text{mis-ID}}-P_{\text{corr-ID}}$
+where $H(S)$ is the average cluster entropy (encouraging homogeneity), $P_{\text{mis-ID}}$ is the proportion of ID samples misclassified as OOD, and $P_{\text{corr-ID}}$ is the proportion of correctly classified ID samples.
+This balances three goals: coherent clusters, low ID misclassification, and high correct ID classification, promoting better surrogate label quality.
 Table 2: Performance metrics comparing the oracle classifier $g_{oracle}$ and the surrogate classifier $g^{*}$ for EuroSAT and xBD datasets across various experimental setups.
@@
 5 Experimental Setup
-In this section, we describe the experimental setups designed to evaluate TARDIS under controlled distribution shifts, which mimic real-world geospatial deployment challenges.
-Specifically, we use the EuroSAT [9] dataset for patch-level classification and the xBD [7] dataset for semantic segmentation.
-EuroSAT contains 27,000 labeled images across ten land use and land cover classes: Annual Crop, Forest, Herbaceous Vegetation, Highway, Industrial, Pasture, Permanent Crop, Residential, River, and Sea/Lake.
-The xBD dataset includes over 1 million labeled samples for building damage assessment from satellite imagery, covering multiple disaster types such as hurricanes, floods, and wildfires.
-For our experiments, we formulate xBD as a building footprint segmentation task with two classes (building footprint and background).
-EuroSAT and xBD provide a suitable test-bed for our OOD method due to their differences in volume, task, sensor type, geospatial distribution, and temporal and spatial characteristics, allowing a thorough evaluation under conditions commonly encountered in geospatial deployments.
-To introduce controlled distribution shifts, we reformulate the train, validation, and test splits of EuroSAT and xBD so that the model, when trained on the modified training set and evaluated on the adjusted test set, encounters shifts similar to those seen in geospatial model deployments.
-We categorize these shifts as covariate shift and semantic shift.
-Covariate shift occurs when the input data distribution changes between training and testing.
-This can introduce selection bias, as covariates in the training set may still appear in the test set to some extent, potentially skewing robustness evaluation.
-To address this, we also introduce semantic shift experiments where a class is held out during testing, creating an unseen semantic class.
-For our experiments, covariate shifts include spatial variance (varying spatial locations), temporal variance (varying time of data collection), and environmental variance (varying landcover or disaster types).
-For semantic shifts in the EuroSAT dataset, we train the model on nine out of ten classes and test on the holdout class, repeating this process so that each class serves as the holdout class once.
-Details of the experimental setups are provided in Table 1.
-The setups described above assume clear definitions of ID (training set) and OOD (test set), which presupposes access to test-time distribution shifts, an assumption that is not feasible in real-world deployments.
-To evaluate our method’s performance under unknown test-time distribution shifts, we fit and evaluate the binary classifier $g$ in two distinct ways.
-In the first approach, we bypass surrogate label assignment (Step 3 in Figure 2), using the well-defined ID and OOD samples to fit and evaluate the classifier, referred to as $g_{\text{oracle}}$.
-In the second approach, we treat the OOD labels in the test set as unknown, simulating real-world conditions where these labels are unavailable.
-This method uses surrogate label assignment to reassign unknown labels into surrogate ID and surrogate OOD.
-Parameter estimation and evaluation on these reassigned labels yield a classifier denoted as $g^{*}$, representing our proposed method.
-The underlying assumption is that $g_{\text{oracle}}$ provides an upper bound on performance for $g^{*}$, as $g_{\text{oracle}}$ is trained with clearly separated ID and OOD samples, whereas $g^{*}$ is trained on surrogate ID and surrogate OOD samples that are generated through our surrogate label assignment process.
-Figure 3: Geographical distribution of ID and WILD sets, containing 500 and 1200 samples, respectively.
-The ID set is sampled from the FTW dataset training set, while the WILD set is randomly sampled from the Microsoft Planetary Computer.
-Each Sentinel-2 patch, provided in two different time frames (planting and harvesting).
-The model $f$ takes both Sentinel-2 images from different seasons as input, and the predictions are shown on the right.
-The bar indicates the probability output of $g$ for each input pair, with a higher probability (closer to 1) suggesting a higher likelihood of the pair being OOD.
-On the map, probability values are thresholded at 0.5.
+In this section, we describe experimental setups to evaluate TARDIS under controlled distribution shifts, mimicking real-world geospatial deployment challenges.
+We use the EuroSAT [12] dataset for patch-level classification and the xBD [10] dataset for semantic segmentation.
+EuroSAT contains 27,000 labeled images across ten land use/land cover classes.
+xBD includes pre/post-disaster imagery labeled for building damage assessment, covering hurricanes, floods, wildfires, and more.
+For our purposes, we formulate xBD as a building footprint segmentation task with two classes: building footprint and background.
+These datasets offer diverse evaluation conditions in terms of volume, task, sensor, spatial/temporal coverage, and geolocation, making them suitable test-beds for our OOD method.
+To simulate controlled distribution shifts, we adjust the train, validation, and test splits in both datasets so that the model, trained on the modified training set, encounters realistic shifts during testing.
+We evaluate two shift types: covariate shift, where input distribution changes between train and test, and semantic shift, where unseen classes appear at test time.
+Covariate shifts include spatial (different locations), temporal (different time periods), and environmental (different landcover/disaster types) variance.
+For semantic shift, we train on nine of ten classes and test on the held-out class, rotating through all classes.
+Table 1 summarizes these setups.
+These setups assume access to clear ID (train) and OOD (test) labels—an unrealistic assumption in real-world deployments.
+To address this, we evaluate our method in two ways.
+First, we use known ID and OOD labels to directly train and evaluate the binary classifier $g$, referred to as $g_{\text{oracle}}$.
+Second, we treat test-time OOD labels as unknown (i.e., WILD) and apply our surrogate label assignment to generate surrogate ID and OOD labels.
+This version, trained and evaluated using surrogate labels, is denoted as $g^{*}$.
+We consider $g_{\text{oracle}}$ as an upper bound on performance, since it uses ground-truth labels, while $g^{*}$ relies on inferred surrogates.
 6 Experimental Results
-In this section, we present the ablation studies and discuss the results of the surrogate sample assignment benchmark, where we evaluate the performance of $g^{*}$.
-Note that ID task performance is not reported, as it remains unaffected by our method.
+In this section, we present ablation studies and experimental results. Note that ID task performance is not reported, as it remains unaffected by our method.
 Ablation Studies.
@@
 This ability depends significantly on the layer from which internal activations are extracted and the downsampling method applied.
-To address this, we test different layers and compare their performance.
+To address this, we test different layers and compare their performance (Table 4 and Table 5).
 For downsampling, we experiment with several methods, including mean and standard deviations, average pooling, max pooling, and PCA-based reduction.
-The max pooling-based downsampling method achieves the highest performance, likely due to its ability to retain the most salient activation patterns, which we argue is important for effective OOD detection.
-After selecting the optimal layer and downsampling method, we evaluate various classifiers: KNeighbors, GaussianNB, DecisionTree, ExtraTrees, LogisticRegression, SVC, RandomForestUnbalanced, RandomForest, AdaBoost, and GradientBoosting.
-Results indicate that Logistic Regression provides the best tradeoff between classification performance and wall time.
+The max pooling-based downsampling method achieves the highest performance, likely due to its ability to retain the most salient activation patterns, which we argue is important for effective OOD detection (Table 6).
+We then evaluate various classifiers: KNeighbors, GaussianNB, DecisionTree, ExtraTrees, LogisticRegression, SVC, RandomForestUnbalanced, RandomForest, AdaBoost, and GradientBoosting.
+Results indicate that Logistic Regression provides the best tradeoff between classification performance and wall time (Table 7).
 We then compare $g^{*}$ to $g_{\text{oracle}}$ as we tune the parameters $k$ and $T$ required for surrogate label assignment.
@@
 Fixing $T$ based on the hyperparameter search results, we observe a recurring pattern that enables us to fix $k$ to 0.3 $\times$ M across all setups.
-This choice is driven by the observation that the classifier $g^{*}$ reaches performance levels close to $g_{\text{oracle}}$ when $K$ is set to 30% of the total training samples.
-This trend holds across both datasets, with the xBD disaster scenario (semantic shift) and EuroSAT Pasture experiment (covariate shift) showing similar behavior.
-In both cases, classifier performance improves as more clusters are incorporated, approaching the performance of $g_{\text{oracle}}$.
-We present the results of TARDIS across these controlled distribution shifts in the Supplementary Material.
+This choice is based on the observation that the classifier $g^{*}$ reaches performance levels close to $g_{\text{oracle}}$ when $k$ is set to 30% of the training samples.
+This trend holds for both datasets, with the xBD disaster scenario (semantic shift) and the EuroSAT Pasture experiment (covariate shift) showing similar behavior.
+In both cases, performance improves as more clusters are added, approaching that of $g_{\text{oracle}}$.
+In Figures 8 and 9, we show how we determine the optimal $(k,T)$ pair that achieves high surrogate label assignment performance across all setups.
+The same pair generalizes to real-world deployment, removing the need for hyperparameter tuning.
 Surrogate Sample Assignment Benchmark.
-The quantitative results for $g^{*}$ and $g_{\text{oracle}}$ across 16 experimental setups are shown in Table 2.
-Across all setups, $g^{*}$ closely matches the performance of $g_{\text{oracle}}$, which represents the upper bound for our method.
-Specifically, in 13 setups for AUROC and 17 setups for FPR95, the difference between $g^{*}$ and $g_{\text{oracle}}$ is not statistically significant.
-In the remaining setups, while a small performance difference exists, the proposed method largely recovers the theoretical upper bound.
-This minimal, and often statistically insignificant, performance gap demonstrates that TARDIS can assign surrogate ID and OOD labels with high performance, even without access to test-time distribution during training.
-A key takeaway from these experiments is the method’s ability to disentangle WILD samples with unknown test-time distributions into surrogate ID and OOD samples.
-Notably, under unknown distribution settings, it achieves performance comparable to cases with known ID and OOD labels, highlighting its strength in reliable OOD detection without prior distribution knowledge, making it suitable for real-world geospatial applications.
+Table 2 shows results for $g^{*}$ and $g_{\text{oracle}}$ across 17 setups.
+Overall, $g^{*}$ closely approaches $g_{\text{oracle}}$, which serves as the upper bound.
+In 13 setups for AUROC and all 17 for FPR95, the performance gap is not statistically significant.
+In the few remaining cases, the gap is small.
+These results show that TARDIS can effectively assign surrogate labels without access to test-time labels.
+ID/OOD Balance.
+As discussed in Section 4, splitting ID and WILD data into validation and clustering sets affects the ID/OOD ratio throughout the pipeline.
+In Table 2, WILD consists entirely of OOD samples, but in real deployments, it may contain both ID and OOD.
+While data is shuffled, monitoring this ratio is key to understanding sensitivity.
+For example, in the Pasture semantic shift setup, the clustering and classifier stages see 17% and 15% OOD, respectively.
+In Nepal Flood – Midwest Flood, the ratios are 46% and 35%.
+This variation shows that TARDIS performs robustly across different ID/OOD balances, supporting the generalizability of our method.
+Figure 3: Geographical distribution of ID and WILD sets, containing 500 and 1200 samples, respectively.
+The ID set is sampled from the FTW dataset training set, while the WILD set is randomly sampled from the Microsoft Planetary Computer.
+Each Sentinel-2 patch, provided in two different time frames (planting and harvesting).
+The model $f$ takes both Sentinel-2 images from different seasons as input, and the predictions are shown on the right.
+The OOD probability values are thresholded at 0.5.
+Table 3: Performance comparison of post-hoc score-based and activation-based OOD detection methods across 10 semantic shift and 7 covariate shift experiments on EuroSAT and xBD.
+We report the mean ± standard deviation across all 10 or 7 experiments for each method.
+TARDIS achieves comparable performance while requiring no hyperparameter tuning and preserving main task performance.
+[…]
+Semantic Shift Mahalanobis [19] 0.85 $\pm$ 0.01 0.19 $\pm$ 0.05
+MSP [13] 0.90 $\pm$ 0.03 0.23 $\pm$ 0.07
+Energy [21] 0.89 $\pm$ 0.02 0.20 $\pm$ 0.04
+ReAct [30] 0.92 $\pm$ 0.04 0.13 $\pm$ 0.01
+NAP [25] 0.93 $\pm$ 0.03 0.21 $\pm$ 0.05
+TARDIS (Ours) 0.95 $\pm$ 0.02 0.16 $\pm$ 0.03
+Covariate Shift Mahalanobis [19] 0.88 $\pm$ 0.03 0.22 $\pm$ 0.06
+MSP [13] 0.92 $\pm$ 0.04 0.25 $\pm$ 0.08
+Energy [21] 0.91 $\pm$ 0.03 0.22 $\pm$ 0.05
+ReAct [30] 0.94 $\pm$ 0.05 0.15 $\pm$ 0.02
+NAP [25] 0.95 $\pm$ 0.02 0.23 $\pm$ 0.06
+TARDIS (Ours) 0.98 $\pm$ 0.01 0.11 $\pm$ 0.04
+Comparison with Existing OOD Detection Methods.
+We benchmark existing post-hoc activation-based and score-based methods across all semantic and covariate shift experiments, reporting mean and standard deviation (Table 3).
+Gradient-based methods involving backpropagation are excluded due to their high computational cost and increased wall time.
+Score-based methods (Mahalanobis, MSP, Energy) preserve ID performance as they require no model changes, but rely on potentially overconfident predictions and are less sensitive to near-domain shifts.
+Activation-based methods like ReAct and NAP modify internal activations—via clipping or purification—boosting OOD detection but at the cost of reduced ID performance, limiting their practicality.
+Why Clustering Alone is Insufficient.
+Our method includes a distance-based clustering step followed by an additional classification step (Step 3 and 4 in Figure 2).
+Removing the classification step results in a drop in ROC AUC for the River experiment from 97% to 79%—a trend consistently observed across all experiments.
+This demonstrates that clustering alone (i.e., distance-based assessment) is insufficient for effective distribution detection, justifying the need for Step 4 in Figure 2.
 7 OOD Detection Goes Global: Real-World Deployment
-We use the Fields of the World (FTW) pre-trained models [11] to demonstrate the capabilities of TARDIS in a large-scale deployment scenario.
-FTW is a geographically diverse dataset designed for agricultural field segmentation, covering 24 regions across Europe, Africa, Asia, and South America: Austria, Belgium, Brazil, Cambodia, Corsica, Croatia, Denmark, Estonia, Finland, France, Germany, India, Kenya, Latvia, Lithuania, Luxembourg, Netherlands, Portugal, Rwanda, Slovakia, Slovenia, South Africa, Spain, Sweden, and Vietnam.
+We use the Fields of the World (FTW)  [14] training set and pre-trained models to demonstrate the capabilities of TARDIS in a large-scale deployment scenario.
+FTW is a geographically diverse dataset designed for agricultural field segmentation, covering 24 regions across Europe, Africa, Asia, and South America.
 The dataset contains approximately 70,000 samples, each consisting of multi-date, multi-spectral Sentinel-2 satellite patches paired with three-class semantic segmentation masks (field, field boundary, and background).
 The task involves segmenting these classes using a pair of Sentinel-2 images — one for the planting season and one for the harvesting season — as input.
-Field boundary data is crucial for global agricultural monitoring, however training large scale models to segment field boundaries from satellite imagery presents significant challenges due to the geographic diversity of fields, varying crop cycles, and agro-climatic conditions, all of which which introduce substantial distribution shifts.
+Field boundary data is crucial for global agricultural monitoring, however training large scale models to segment field boundaries from satellite imagery presents significant challenges due to the geographic diversity of fields, varying crop cycles, and agro-climatic conditions, all of which introduce substantial distribution shifts.
 This complexity, along with corresponding distribution shifts found in real world imagery, makes the FTW dataset ideal for testing TARDIS’s ability to diagnose model performance during inference.
-Additionally, the multi-date nature of the dataset is particularly suitable for evaluating segmentation models that must handle spatiotemporal variations in satellite imagery.
+Additionally, the multi-date nature of the dataset is particularly suitable for evaluating models that must handle spatiotemporal variations in satellite imagery.
 Sampling ID Set.
@@
 Sampling WILD Set.
-We collect multispectral Sentinel-2 satellite images using the Microsoft Planetary Computer [22], which is freely accessible.
+We collect multispectral Sentinel-2 L2A satellite images using the Microsoft Planetary Computer [28].
 The images are processed to Level-2A (bottom-of-atmosphere) and stored in cloud-optimized GeoTIFF (COG) format.
@@
 The two patches are concatenated along the channel dimension, resulting in a patch of shape 256 $\times$ 256 $\times$ 8 pixels.
-Following the pre-trained model’s convention, the concatenated patch is upsampled by a factor of 2 to 512 $\times$ 512 $\times$ 8, processed by the model, and then downsampled back to the original dimensions.
+Following the pre-trained model’s convention, the concatenated patch is upsampled to 512 $\times$ 512 $\times$ 8, processed by the model, and then downsampled back to the original dimensions.
 For surrogate label assignment, we fix the hyperparameters $k$ and $T$ to $0.3\times 1200$ and 0.1, respectively, as suggested by the benchmark study.
-The geographical distribution of ID and WILD samples is illustrated in Figure 3, where WILD samples are further classified into ID and OOD categories.
-A notable pattern is that samples located in arid biomes — such as the deserts in Inner Australia, the Sahara, and Patagonia — and polar regions, including Icelandic glaciers and the South Pole, are more likely to be assigned as OOD.
-This can be attributed to the ID samples predominantly representing mesic environments, which are areas with moderate moisture levels and managed landscapes suitable for agriculture.
-In contrast, arid biomes and polar regions exhibit extreme environmental conditions that differ markedly from the mesic environments of the ID samples.
-This ecological dissimilarity results in these biomes being classified as surrogate OOD, highlighting the model’s sensitivity to environmental context and distribution shifts.
-To evaluate the scalability of our method, we measure the time required to classify the internal activation of a WILD sample as surrogate ID or OOD using the classifier $g$, implemented as a logistic regressor.
-An input patch of shape $256\times 256$ pixels, with a spatial resolution of 10 meters per pixel, covers an area of $6.5536\,\text{km}^{2}$ and takes 0.003 seconds to process.
-Applied to the entirety of mainland Africa, which spans $29.77\times 10^{6}\,\text{km}^{2}$, the method could be deployed on a single CPU in approximately 15 days.
-In a distributed GPU setup, this process could be completed within hours (a fraction of the time required for $f$ model inference over a similar area).
-This is crucial as such a model would allow practitioners to see, spatially, where the field boundary model itself is performing in distribution versus not (and where the corresponding outputs are likely irrelevant).
-These results demonstrate the practicality of the proposed method for deployment in low-data regimes, where geospatial models are typically less robust, and highlight its potential for improving model understanding in these challenging settings.
-We further evaluate the reliability of the OOD classifier $g$ on the training set by comparing the performance of the $f$ model on the FTW dataset’s test set with the probability scores from $g$.
-We observe a correlation between the performance of the $f$ model on the test set and low skewness in the output of $g$.
-This indicates that when the $f$ model performs poorly, the OOD classifier $g$ reflects this by producing unimodal OOD scores, as shown by low skewness values (see Figure 4) for countries such as Portugal, Cambodia, and Vietnam.
-Figure 4: Relationship between model performance and OOD score skewness.
-The figure illustrates the correlation between low performance of the $f$ model on the FTW test set and low skewness in the OOD classifier $g$’s score distribution.
-When the $f$ model performs poorly, the OOD classifier $g$ produces more OOD scores, as reflected by lower skewness values.
+The geographical distribution of ID and WILD samples is shown in Figure 3, with WILD samples further classified into surrogate ID and OOD.
+A clear pattern emerges: samples from arid biomes—such as the deserts of Inner Australia, the Sahara, and Patagonia—and polar regions, including Icelandic glaciers and the South Pole, are more frequently assigned as OOD.
+This is likely because the ID samples predominantly represent mesic environments: moderately moist, managed landscapes typical of agricultural areas.
+In contrast, arid and polar regions exhibit extreme environmental conditions that differ significantly from those in the ID set.
+This ecological dissimilarity likely drives their classification as surrogate OOD, reflecting the model’s sensitivity to environmental context and distribution shifts.
+To evaluate scalability, we measure the time required to classify a WILD sample’s internal activation as surrogate ID or OOD using the logistic regression classifier $g$.
+A $256\times 256$ input patch (10 m resolution, covering $6.5536\,\text{km}^{2}$) takes 0.003 seconds to process.
+Applied to mainland Africa ($29.77\times 10^{6}\,\text{km}^{2}$), the method runs on a single GPU in under 4 hours—a fraction of the time needed for full $f$-model inference.
+This enables spatial diagnostics, identifying where the field boundary model is likely in- or out-of-distribution and where outputs may be unreliable.
+Such rapid screening is especially useful in low-data regions, where models are typically less robust and OOD detection can guide targeted interventions or data collection.
+We assess the reliability of the OOD classifier $g$ by comparing its output skewness on the training set to the performance of the model $f$ on the FTW test set.
+High skewness indicates confident (mostly ID or OOD) predictions, while low skewness suggests uncertainty or misclassification.
+Plotting skewness against $f$’s test performance reveals a trend: low skewness aligns with poor model performance (Figure 4), as seen in countries like Portugal, Cambodia, and Vietnam.
+This suggests that low skewness may signal the presence of OOD samples.
+Thus, TARDIS can also serve as a sanity check for detecting distribution shifts in test splits.
+Limitations and Future Work.
+Our method assumes access to a non-trivial number of WILD samples and requires periodic updates to the clustering used for surrogate label assignment.
+Future work could explore density-based clustering methods instead of the current distance-based approach, as they may better handle variable sample densities and outliers, further improving surrogate label assignment.
+Figure 4: Correlation between model performance and OOD score skewness.
+Low performance of the $f$ model on the FTW test set aligns with low skewness in the OOD classifier $g$’s score distribution, suggesting the presence of OOD samples in the test set.
 8 Conclusion
-We present TARDIS, which effectively assigns surrogate ID or OOD labels to samples with unknown distributions by clustering the representation space.
-These surrogate labels are used to train a binary classifier for OOD detection.
-We do not aim to optimize OOD detection performance but instead propose a method designed for real-world applications where the distribution is unknown, maintaining ID task performance is critical, and computationally expensive methods are impractical.
-It achieves performance close to the theoretical upper limit in assigning surrogate ID and OOD samples in 13 out of 17 experiments.
-We demonstrate TARDIS can be deployed at scale, with the potential to enhance the robustness and trustworthiness of geospatial models as well as provide insights into distribution shifts in global applications.
+We present TARDIS, a distribution shift detector that assigns surrogate ID and OOD labels to samples from unknown distributions by leveraging clustering in the feature space.
+These surrogate labels are assigned based on known ID samples and used to train a binary classifier that outputs a distribution score during inference.
+Rather than optimizing for OOD detection performance, TARDIS targets real-world scenarios where the distribution is unknown, ID task performance must be preserved, and computational overhead is a concern.
+Across 13 of 17 experiments mimicking real-world semantic and covariate distribution shifts, our method achieves near-upper-bound accuracy for surrogate label assignment while matching the performance of top post-hoc OOD detection methods.
+We demonstrate how TARDIS scales in a real-world application, providing interpretable insights into model behavior under distribution shifts and helping identify potential biases and limitations in the dataset.
+This contributes to the robustness and trustworthiness of models, making it particularly valuable in safety-critical, time-sensitive, and low-data settings.
 Supplementary Material
@@
 We simplify the damage assessment task into binary segmentation by reassigning damage levels: background (0) and levels 1-2 are grouped, while levels 3-4 form a high-damage class.
-This minimizes concept drift and ensures a balanced evaluation.
+This minimizes concept drift and ensures a fair evaluation of distribution shifts.
 We train a U-Net model with a ResNet50 backbone, pre-trained on ImageNet and configured for 3 input channels.
@@
 These samples demonstrate the temporal and semantic differences between pre- and post-disaster scenes, highlighting the challenges posed by distribution shifts.
-Figure 6: Examples from the dataset, with one sample from each class. These images highlight the spatial and semantic distinctions across classes.
-Figure 7: The input samples are from the WILD set, where the distribution is unknown.
+Figure 6: Examples from the EuroSAT dataset, with one sample from each class. These images highlight the spatial and semantic distinctions across classes.
+Figure 7: Deploying TARDIS over FTW dataset: The input samples are from the collected WILD set, where the distribution is unknown.
 The figure shows Sentinel-2 images at two different times (planting season and harvesting season — Window A and Window B).
@@
   4. 4.
-    PCA-based reduction (PCA): Applies Principal Component Analysis (PCA) to reshape the activation map into a vector and projects it into a lower-dimensional space with 10 components.
+    PCA-based reduction (PCA): Applies Principal Component Analysis to reshape the activation map into a vector and projects it into a lower-dimensional space with 10 components.
 We summarize the OOD detection performance across all experiments on the EuroSAT and xBD datasets under different downsampling methods in Table 6, using the FPR95 metric.
@@
 The underlying assumption is that samples with similar distributions lie closer in the activation space than those from dissimilar distributions.
-Effectively segmenting the activation space is critical, as the distributions of WILD samples are unknown during deployment.
-Segmenting nearby activations and correctly assigning clusters to surrogate ID or OOD makes the choice of k and T crucial for accurate OOD detection.
+Effectively clustering the activation space is critical, as the distributions of WILD samples are unknown during deployment, and it depends on the optimal choice of $k$ and $T$.
 To develop insights into selecting k and T, we conduct controlled experiments on EuroSAT and xBD, where ID and OOD labels are known.
@@
 The primary goal is to understand how to choose k and T, and whether there are patterns we can extrapolate to real-life deployment.
-For this, we calculate the ratio of k to the total number of training samples and evaluate its effect on OOD detection performance (Accuracy, FPR95, and AUROC).
+For this, we first assign surrogate labels and calculate the ratio of k to the total number of training samples and evaluate its effect on OOD detection performance (Accuracy, FPR95, and AUROC).
 We plot these metrics against the ratio of k/total training samples, increasing k until the ratio reaches 1.
-Theoretically, OOD detection improves with more clusters as this enables finer-grained segmentation of the activation space, reducing the risk of including anomalies in ID clusters.
-To establish a theoretical maximum (oracle performance), we also evaluate OOD classification with known ID and OOD labels, bypassing the need for clustering.
-This oracle performance is represented by horizontal dashed lines in Figures 8 and 9 (upper plots).
+Theoretically, OOD detection improves with more clusters as this enables finer-grained clustering of the activation space, reducing the risk of including anomalies in ID clusters.
+To establish a theoretical maximum (upper-bound performance), we also evaluate OOD classification with known ID and OOD labels, bypassing the need for clustering.
+This oracle performance is represented by horizontal dashed lines in Figure 8 and Figure 9 (upper plots).
 The results for two representative experiments—one from EuroSAT and one from xBD—since all experiments show similar trends.
@@
 The absence of degradation in performance further underscores the importance of activation-level clustering as a reliable proxy for domain estimation based on neighboring samples.
-We set k and T as described and use t-SNE in the lower plots of Figures 8 and 9 to reduce the dimensionality of the activation spaces to 2D for visualization.
+We set k and T as described and use t-SNE in the lower plots of Figure 8 and Figure 9 to reduce the dimensionality of the activation spaces to 2D for visualization.
 When ID and OOD labels are known, the t-SNE plots show that only a small fraction of labels changes from the original labels.
 This demonstrates the effectiveness of the surrogate label assignment process described above.
-11 Additional Experimental Results
+11 Further Experimental Results
 In Figure 10, we show the predictions of the DL model $f$ and the OOD classifier $g$, along with the ground truth class and distribution annotations for the EuroSAT experiment, where Forest serves as the OOD class.
@@
 The performance of $f$ on the test set is not directly measurable since the test uses a single unseen class.
-We report the performance of $g$ as: Accuracy = 0.9325, ROC AUC = 0.9886, and FPR95 = 0.0619.
-For xBD, we present results where $f$ trains on Hurricane Matthew (ID, Figure 11) and tests on Mexico Earthquake (OOD, Figure 12).
-Comparing the input images and masks between ID and OOD shows that even when $f$ performs suboptimally, $g$ effectively quantifies the distribution shifts.
-The performance of $f$ on the test set is: Multiclass Accuracy = 0.7690 and Multiclass Jaccard Index = 0.6248.
-We attribute $f$’s suboptimal prediction performance to the significant distribution shift between the training (Hurricane Matthew) and testing (Mexico Earthquake) datasets.
-The performance of $g$ is: Accuracy = 0.9806, ROC AUC = 0.9986, and FPR95 = 0.0.
+We report the performance of $g$ as: Accuracy: $93.25\%$, ROC AUC: $98.86\%$, FPR95: $6.19\%$.
+For xBD, we present results where $f$ is trained on Hurricane Matthew (ID, Figure 11) and tested on Mexico Earthquake (OOD, Figure 12).
+Comparing the input images and masks between ID and OOD reveals that even when $f$ performs suboptimally, $g$ effectively quantifies the distribution shifts.
+The performance of $f$ on the test set is as follows: Multi-class accuracy: $76.90\%$, Multi-class Jaccard index: $62.48\%$.
+We attribute $f$’s suboptimal prediction performance to the significant distribution shift between the training (Hurricane Matthew) and testing (Mexico Earthquake) datasets, and also to the fact that we reformulate the main task of damage classification to building detection (as described in Section 9.2).
+The performance of $g$ is: Accuracy: $98.06\%$, ROC AUC: $99.86\%$, FPR95: $0.00\%$.
 Experiment 2/217 8/217 16/217 38/217 43/217 48/217 118/217 139/217 199/217 211/217
@@
 GradientBoosting 92.92 ± 0.81 93.11 ± 0.85 30.47 ± 1.88 7.00 ± 2.00
-Table 6: Benchmark results of classifiers $g$, including Accuracy, ROC_AUC, FPR95, and prediction time.
+Table 7: Benchmark results of classifiers $g$, including Accuracy, ROC AUC, FPR95, and prediction time.
 Values are reported as mean ± SEM over all experiments on EuroSAT and xBD.